OpenStudy (amistre64):

a function is linear if: f(a+b) = f(a) + f(b) Let f(x) = 2x+3 ; is this linear?

OpenStudy (amistre64):

im curious if we would have to move this to pass thru the origin for this definition of work out

OpenStudy (amistre64):

does the constant at the end matter?

OpenStudy (raden):

f(x) = 2x+3 f(1) = 2(1)+3 = 5 f(2) = 2(2)+3 = 7 f(3) = 2(3) + 3 = 9 f(3) = f(1+2) = f(1) + f(2) = 5 + 7 = 12 nah, this is not same like the defenition above :)

OpenStudy (anonymous):

But @RadEn u cant generalize

OpenStudy (amistre64):

im thinking that since it doesnt contain the point (0,0); since its of the family 2x+ k,

OpenStudy (anonymous):

f(x)+f(x+3)=4x+12, f(2x+3)=4x+9

OpenStudy (amistre64):

2(a+b) = 2(a) + 2(b)

OpenStudy (amistre64):

im not sure of its proper to say that {2x+k} is a congruence class or not

OpenStudy (anonymous):

congruence class??

OpenStudy (anonymous):

Congruence class? Does it have anything to do with the linearity of a function?

OpenStudy (amistre64):

a congruence class is a set of elements that are by all rights ... equaivalent like 7 mod 6 = 13 mod 6 because 7 and 13 are both equal to 1 mod 6 they are in the same congruence class mod 6

OpenStudy (amistre64):

if all functions of the form: {2x + k} are in the same congruence class, then it would most likely be proper to say that f(x) = 2x+3 is linear, since f(x) = 2x+0 is linear

OpenStudy (anonymous):

Why not just use the defining property of a linear function, which was already given in the problem?

OpenStudy (amistre64):

becasue: 2(a+b) + 3 is not equal to 2a + 3 + 2b + 3

OpenStudy (amistre64):

but ... since the line 2x+3 and 2x are parallel, wouldnt the properties of 2x be the same as 2x+3?

OpenStudy (amistre64):

barring of course the y intercept

OpenStudy (anonymous):

Yes, I have the same concern, but if the definition says it's not linear, then it's not.

OpenStudy (anonymous):

I think the vector space section explains it pretty well: http://en.wikipedia.org/wiki/Linear_function

OpenStudy (amistre64):

but im sure that definition is based on .. i forget the name, kernel maybe? ... such that 2x+3 is not a proper vector space (or someting) since it doesnt contain the (0,0) point

OpenStudy (amistre64):

pfft, i can never read those wiki with any degree of sensibility :)

OpenStudy (anonymous):

Then go search linear function vs. affine function. Basically my intuition was just wrong, a function f(x)=ax+b is called a linear function iff b=0, though its graph is still line-ar.

OpenStudy (amistre64):

i thought about googling it later today, just thought id bring it up in here for people to be able to ponder something other than ... well, most of whats posted here

OpenStudy (amistre64):

affine is something i need to get better acquainted with fer sure

OpenStudy (anonymous):

Consider posting it up on Math SE if you want a better insight into it.